# Selective Population Protocols

### v2

@@ Abstract
The model of population protocols provides a universal platform to study distributed processes driven by pairwise interactions of anonymous agents. Pairs of agents are drawn by the random scheduler uniformly at random from a large population of size $n$. Each interacting pair is formed of the initiator and the responder. On the conclusion of an interaction the states of the two agents are modified according to the predefined transition function, i.e., the set of rules underpinning the considered process. The state space $S$ of agents is fixed and the size of the population is not known, i.e., not hard-coded in the transition function. The time performance of a protocol is measured by the number of interactions $|{\mathcal{I}}|$ required to reach a final configuration, or expressed in terms of parallel time (or simply time) defined as $|{\mathcal{I}}|/n.$ We say that a protocol is efficient if it stabilises in parallel time $O(\mbox{poly}\log n)$. Finally, we say that a population protocol is stable if it stabilises in a final configuration asserting the correct solution (with probability 1).
While population protocols present an elegant and robust model for randomized distributed computation, their efficiency wanes when tackling issues that require more focused communication or the execution of multiple processes. This is a result of their inherent dependence on global communication. To address this issue, we propose a new, selective variant of population protocols by introducing a partition of the state space and the corresponding conditional selection of responders. We demonstrate on several examples that the new model offers a natural environment, complete with tools and a high-level description, to facilitate more efficient solutions.
In particular, we provide fixed-state stable and efficient solutions to two central problems: leader election and majority computation, both with confirmation. This constitutes a separation result, as achieving stable and efficient majority computation requires $\Omega(\log n)$ states in standard population protocols [2], even when the leader is already determined. Additionally, we explore the computation of the median using the comparison model, where the operational state space of agents is fixed, and the transition function determines the order between (arbitrarily large) hidden keys associated with interacting agents. Our findings reveal that the computation of the median of $n$ numbers requires $\Omega(n)$ time. Moreover, we demonstrate that the problem can be solved in $O(n\log n)$ time, both in expectation and with high probability, in standard population protocols. In contrast, we establish that a feasible solution in selective population protocols can be achieved in $O(\log^{4}n)$ time.
@@ 2 Selective Population Protocols
One-way epidemic Consider a communication primitive known as one-way epidemic [6] in which state $1$ of the source agent is propagated to all other agents initially being in state $0$ The transition function has only one rule in the standard population protocol model
$1+0\rightarrow 1+1$
@@
$(1)\;$ $0\displaystyle 1+\mathcal{G}_{0}|$ 0 $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $1+1$ $(2)\;$ $1+\mathcal{G}_{0}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $Stop$
@@ 2.1.1 Example of Leader Driven Selective Protocol
$(1)\;$ $L_{in}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{Y}+\text{Free}$ $(2)\;$ $L_{in}+\mathcal{G}_{X}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{out}$ $(3)\;$ $L_{Y}+\mathcal{G}_{Y}|Y$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{Z}+\hat{Y}$ $(4)\;$ $L_{Y}+\mathcal{G}_{Y}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{\hat{Y}}$ $(5)\;$ $L_{Z}+\mathcal{G}_{R}|\text{Free}$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{Y}+Z$ $(6)\;$ $L_{\hat{Y}}+\mathcal{G}_{\hat{Y}}|\hat{Y}$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{\hat{Y}}+Y$ $(7)\;$ $L_{\hat{Y}}+\mathcal{G}_{\hat{Y}}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{in}$
@@ 2.2.1 Fragmented Parallel Time
$(1)\;$ $1\displaystyle 0+\mathcal{G}_{1}|$ 1 $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $1\displaystyle 1^{*}+$ 1 $(2)\;$ $1+\mathcal{G}_{0}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $1^{*}$
If in the initial configuration there is exactly one informed agent in state $1,$ all other agents in state 00 contact this agent to get informed and rest, see rule (1). In the last meaningful interaction rule (2) rests the unique informer. While the number of interactions before stabilisation with all agents resting in state $1^{*}$ is $O(n\log n)$ whp, the parallelism of this epidemic process is very poor as only one agent informs others as the responder.
@@
Let us first estimate the probability that a given chunk corresponds to time smaller than $\ln n$. This probability is not greater than the probability that in time $\ln n$ (starting at the beginning of the chunk) some agent has the responder type interactions greater than $10\ln n$. By Chernoff bound444 We utilise the Chernoff bound variant: $\Pr\left(X>(1+\delta)\mathbb{E}X\right)<\exp\left(-\delta^{2}\mathbb{E}X/(1+\delta)\right)$ for $\delta>0$. , the probability that in time $\ln n$ a given agent experiences $X>10\ln n=10\mathbb{E}X$ interactions as the responder can be estimated by
@@
Let us divide all interactions into maximal subseries such that for any fixed agent there are at most $110\ln n-$ 1 events $A$ involving this agent. Note that these subseries are simultaneously chunks of interactions in which any agent acts as the responder at most $10\ln n$ times since any agent is a responder in $I$ in an external interaction at most once. Let us first estimate the probability that a given subseries has less than $n\ln n$ interactions. This probability is not greater than the probability that during $n\ln n$ interactions (counting from the beginning of the subseries) event $A$ happens for some agent at least $10\ln n$ times. Using calculations from Lemma 1, one can estimate that this probability for a specific agent is smaller than $n^{-81/11}$.
@@
Let us divide interactions into maximal subseries such that for any fixed agent there are at most $110\ln n-$ 1 events $A$ involving this agent. Note that these subseries are simultaneously chunks of interactions in which any agent acts as the responder at most $10\ln n$ times since any agent is a responder in $|I|$ at most once in an external interaction with a non-leader. Let us first estimate the probability that a given subseries has less than $n\ln n$ interactions. This probability is not greater than the probability that during $n\ln n$ interactions (starting at the beginning of the subseries) event $A$ happens for some agent at least $10\ln n$ times. By Chernoff bound, the probability that for a given agent event $A$ in $n\ln n$ interactions occurs $X\geq 10\ln n\geq 5\mathbb{E}X$ can be estimated as follows
@@
By the union bound the probability that in a given subseries event $A$ occurs for any agent more than $110\ln n-$ 1 times does not exceed $n^{-13/3}$. Thus, for $n$ large enough and $k>\frac{|I|}{10n\log n}>\frac{6}{13}\eta$ the probability that at least half of $k$ subseries correspond to parallel time smaller than $\ln n$ is less than
@@ 2.3 Leader Election
$(1)\;$ $L+\mathcal{G}_{0}|L$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L+F$ $(2)\;$ $L+\mathcal{G}_{0}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L^{*}$ $(3)\;$ $L^{*}+\mathcal{G}_{1}|F$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L^{*}+F^{*}$ $(4)\;$ $F^{*}+\mathcal{G}_{1}|F$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $F^{*}+F^{*}$
@@ 2.4 Majority Computation
$(1)\;$ $R+\mathcal{G}_{G}|G$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $N+N$ $(2)\;$ $R+\mathcal{G}_{G}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $R^{*}$ $(3)\;$ $G+\mathcal{G_{R}}|R$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $N+N$ $(4)\;$ $G+\mathcal{G_{R}}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $G^{*}$ $(5)\;$ $R^{*}+\mathcal{G_{N}}|N$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $R^{*}+R^{*}$ $(6)\;$ $G^{*}+\mathcal{G_{N}}|N$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $G^{*}+G^{*}$
@@
In the second phase the number of interactions needed to make all transitions $G\rightarrow G^{*}$ or $R\rightarrow R^{*}$ is $T=T_{k}+T_{k-1}+T_{k-2}+\cdots+T_{1}$, where $T_{i}$ is the number of interactions to reduce the number of $G$ or $R$ agents from $i$ to $i-1$. Note that $T_{i}$ has geometric distribution and $\mathbb{E}T_{i}=n/i$. Thus, by Janson bound $T$ is $O(n\log n)$ whp. Then we count the number of interactions $T^{\prime}$ needed to turn all remaining agents to state $G^{*}$ or $R^{*}$ after initial $T$ interactions of the second phase. We get
@@ 3.1 Median Computation in Standard Model
[…]
@@ 3.2 Fast Median Computation
2:Select randomly leader (as pivot) $p\in S$triangleright leader election
@@
9:    $p\leftarrow$ randomly chosen (by current $p$) agent in $C$triangleright leader hand over
@@
11:return $p$triangleright result announcement
@@ 3.2.1 Partitioning via Coloring
If, after drawing $i-1$ balls, no more than $(i-1)p$ of them are black, we have $q_{i}=\Pr(X_{i}=1|\text{current urn content})\geq p$.
@@
  • If, after drawing $i-1$ balls, more than $(i-1)p$ of them are black, we independently draw $Y_{i}=1$ with a probability of $p$, and $Y_{i}=0$ otherwise.
  • If, after drawing $i-1$ balls, at most $(i-1)p$ of them are black and $X_{i}=0$, then $Y_{i}=0$.
  • If, after drawing $i-1$ balls, at most $(i-1)p$ of them are black and $X_{i}=1$, then $Y_{i}=1$ with a probability of $p/q_{i}$, and $Y_{i}=0$ otherwise.
Variables $Y_{1},Y_{2},\ldots,Y_{\kappa}$ are independent, because no $Y_{i}$ depends on any combination of $Y_{j}:j<i$, and $\Pr(Y_{i})=p$. If $X<(1-\delta)p{\kappa}$, then let $j\in[0,{\kappa})$ be the last draw before which not less than $(j-1)p$ balls drawn are black. This implies $Y_{i}\leq X_{i}$ for $i>j$ and $X_{1}+\cdots+X_{j}\geq p(j-1)$. Thus when $X_{1}+\cdots+X_{\kappa}<(1-\delta)p{\kappa}-p$, we get
@@
We divide the sequence of interactions into interaction periods $[t_{0},t_{1}),[t_{1},t_{2}),[t_{2},t_{3}),\dots$ such that the first interaction period has length $O(n\log^{2}n),$ and each subsequent period is of length $O(n\log n)$. We show that in interaction $t_{\log n}$ all $m$ agents are in the group $\mathcal{G}$ whp. Each colored agent can utilise at most 21 tickets to move other agents to $\mathcal{G}$. This implies that in $t_{\log n}$ at least $m/22$ of these agents have color. Since each interaction between groups removes an agent from ${\mathcal{G}}_{in}$ irreversibly, and there are $O(n\log^{2}n)$ interactions altogether whp, by Lemma 2 the fragmented time of one phase is $O(\log^{2}n)$. The thesis of the lemma is a direct consequence of Lemma 9, which we formulate below. And indeed, since this implies $k(t_{\log n})\geq\min\left\{20\cdot 2^{\log n-1}\log^{2}n,m\right\},$ we get $k(t_{\log n})=m$. ∎
@@
We first deal with the base case $i=1$. Let $t=t_{0}+cn\log^{2}n$, where a constant $c>0$ is to be specified later on, and $t_{1}=t+500n$. Note that the pivot always colors an uncolored agent during an interaction. During interaction period $[t_{0},t)$ the pivot is chosen as the initiator at least $\log^{2}n$ times whp. Indeed, by Chernoff bound the number of times $X,$ for which the pivot is chosen to be an initiator, satisfies for some $c>0$
@@
If in $[t_{0},t)$ group ${\mathcal{G}}_{in}$ runs out of agents, then $k(t)=m$. Otherwise in time $t$ at least $\log^{2}n$ agents are colored and by lemma 7 we have $k(t_{1})=k(t+500n)\geq\min\left\{20\log^{2}n,m\right\}$ whp.
On the other hand, consider an uncolored agent from the first bucket. No matter whether it belongs to ${\mathcal{G}}_{in}$ or ${\mathcal{G}}$, it has probability at least $\frac{1}{nm}$ of interaction with the pivot. Due to the way the buckets are defined, see the proof of Lemma 8, each of them has at least $\frac{m}{4\log n}$ agents. If in any interaction of $[t_{0},t)$ more than $\frac{m}{10\log n}$ agents are colored, then $Inf_{i}(t_{1})>\min\left\{\log n,\frac{m}{10\log n}\right\}$. Otherwise, the probability of coloring a new agent from the first bucket in each interaction of $[t_{0},t)$ is at least $\frac{3}{20n\log n}$. By Chernoff bound, for some $c>0$, the number $X$ of colored agents in the first bucket in $t$ satisfies
@@
then at least $y$ agents from the $i$th bucket are colored in $[t_{i-1},t)$ or in some $\tau$ the number of uncolored agents in $\tau$ is smaller than $\frac{0.9k(\tau)}{8n\log n}$. Both cases imply

### v3

@@ Abstract
The model of population protocols provides a universal platform to study distributed processes driven by pairwise interactions of anonymous agents. While population protocols present an elegant and robust model for randomized distributed computation, their efficiency wanes when tackling issues that require more focused communication or the execution of multiple processes. To address this issue, we propose a new, selective variant of population protocols by introducing a partition of the state space and the corresponding conditional selection of responders. We demonstrate on several examples that the new model offers a natural environment, complete with tools and a high-level description, to facilitate more efficient solutions.
In particular, we provide fixed-state stable and efficient solutions to two central problems: leader election and majority computation, both with confirmation. This constitutes a separation result, as achieving stable and efficient majority computation requires $\Omega(\log n)$ states in standard population protocols, even when the leader is already determined. Additionally, we explore the computation of the median using the comparison model, where the operational state space of agents is fixed, and the transition function determines the order between (arbitrarily large) hidden keys associated with interacting agents. Our findings reveal that the computation of the median of $n$ numbers requires $\Omega(n)$ time. Moreover, we demonstrate that the problem can be solved in $O(n\log n)$ time, both in expectation and with high probability, in standard population protocols. In contrast, we establish that a feasible solution in selective population protocols can be achieved in $O(\log^{4}n)$ time.
@@ 2 Selective Population Protocols
One-way epidemic Consider a communication primitive known as one-way epidemic [6] in which state $1$ of the source agent is propagated to all other agents initially being in state $0.$ The transition function has only one rule in the standard population protocol model
$1+0\rightarrow 1+1.$
@@
$(1)\;$ $1+\mathcal{G}_{0}|0$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $1+1$ $(2)\;$ $1+\mathcal{G}_{0}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $Stop\$
@@ 2.1.1 Example of Leader Driven Selective Protocol
$(1)\;$ $L_{in}+\mathcal{G}_{X}|X$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{Y}+\text{Free}$ $(2)\;$ $L_{in}+\mathcal{G}_{X}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{out}\$$(3)\;$ $L_{Y}+\mathcal{G}_{Y}|Y$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{Z}+\hat{Y}$ $(4)\;$ $L_{Y}+\mathcal{G}_{Y}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{\hat{Y}}\$$(5)\;$ $L_{Z}+\mathcal{G}_{R}|\text{Free}$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{Y}+Z$ $(6)\;$ $L_{\hat{Y}}+\mathcal{G}_{\hat{Y}}|\hat{Y} \displaystyle\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{\hat{Y}}+Y$ $(7)\;$ $L_{\hat{Y}}+\mathcal{G}_{\hat{Y}}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L_{in}\$
@@ 2.2.1 Fragmented Parallel Time
$(1)\;$ $0+\mathcal{G}_{1}|1$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $1^{*}+1$ $(2)\;$ $1+\mathcal{G}_{0}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $1^{*}\$
If in the initial configuration there is exactly one informed agent in state $1,$ all other agents in state 0 contact this agent to get informed and rest, see rule (1). In the last meaningful interaction rule (2) rests the unique informer. While the number of interactions before stabilisation with all agents resting in state $1^{*}$ is $O(n\log n)$ whp, the parallelism of this epidemic process is very poor as only one agent informs others as the responder.
@@
Let us first estimate the probability that a given chunk corresponds to time smaller than $\ln n$. This probability is not greater than the probability that in time $\ln n$ (starting at the beginning of the chunk) some agent has the responder type interactions greater than $10\ln n$. By Chernoff bound [^4] We utilise the Chernoff bound variant: $\Pr\left(X>(1+\delta)\mathbb{E}X\right)<\exp\left(-\delta^{2}\mathbb{E}X/(1+\delta)\right)$ for $\delta>0$. , the probability that in time $\ln n$ a given agent experiences $X>10\ln n=10\mathbb{E}X$ interactions as the responder can be estimated by
@@
Let us divide all interactions into maximal subseries such that for any fixed agent there are at most $10\ln n-1$ events $A$ involving this agent. Note that these subseries are simultaneously chunks of interactions in which any agent acts as the responder at most $10\ln n$ times since any agent is a responder in $I$ in an external interaction at most once. Let us first estimate the probability that a given subseries has less than $n\ln n$ interactions. This probability is not greater than the probability that during $n\ln n$ interactions (counting from the beginning of the subseries) event $A$ happens for some agent at least $10\ln n$ times. Using calculations from Lemma 1, one can estimate that this probability for a specific agent is smaller than $n^{-81/11}$.
@@
Let us divide interactions into maximal subseries such that for any fixed agent there are at most $10\ln n-1$ events $A$ involving this agent. Note that these subseries are simultaneously chunks of interactions in which any agent acts as the responder at most $10\ln n$ times since any agent is a responder in $|I|$ at most once in an external interaction with a non-leader. Let us first estimate the probability that a given subseries has less than $n\ln n$ interactions. This probability is not greater than the probability that during $n\ln n$ interactions (starting at the beginning of the subseries) event $A$ happens for some agent at least $10\ln n$ times. By Chernoff bound, the probability that for a given agent event $A$ in $n\ln n$ interactions occurs $X\geq 10\ln n\geq 5\mathbb{E}X$ can be estimated as follows
@@
By the union bound the probability that in a given subseries event $A$ occurs for any agent more than $10\ln n-1$ times does not exceed $n^{-13/3}$. Thus, for $n$ large enough and $k>\frac{|I|}{10n\log n}>\frac{6}{13}\eta$ the probability that at least half of $k$ subseries correspond to parallel time smaller than $\ln n$ is less than
@@ 2.3 Leader Election
$(1)\;$ $L+\mathcal{G}_{0}|L$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L+F$ $(2)\;$ $L+\mathcal{G}_{0}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L^{*}\$$(3)\;$ $L^{*}+\mathcal{G}_{1}|F$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $L^{*}+F^{*}$ $(4)\;$ $F^{*}+\mathcal{G}_{1}|F$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $F^{*}+F^{*}$
@@ 2.4 Majority Computation
$(1)\;$ $R+\mathcal{G}_{G}|G$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $N+N$ $(2)\;$ $R+\mathcal{G}_{G}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $R^{*}\$$(3)\;$ $G+\mathcal{G_{R}}|R$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $N+N$ $(4)\;$ $G+\mathcal{G_{R}}|null$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $G^{*}\$$(5)\;$ $R^{*}+\mathcal{G_{N}}|N$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $R^{*}+R^{*}$ $(6)\;$ $G^{*}+\mathcal{G_{N}}|N$ $\stackrel{{\scriptstyle}}{{\longrightarrow}}$ $G^{*}+G^{*}$
@@
In the second phase the number of interactions needed to make all transitions $G\rightarrow G^{*}$ or $R\rightarrow R^{*}$ is $T=T_{k}+T_{k-1}+T_{k-2}+\cdots+T_{1}$, where $T_{i}$ is the number of interactions to reduce the number of $G$ or $R$ agents from $i$ to i1i-1. Note that $T_{i}$ has geometric distribution and $\mathbb{E}T_{i}=n/i$. Thus, by Janson bound $T$ is $O(n\log n)$ whp. Then we count the number of interactions $T^{\prime}$ needed to turn all remaining agents to state $G^{*}$ or $R^{*}$ after initial $T$ interactions of the second phase. We get
@@ 3.1 Median Computation in Standard Model
$(1)\;$ $N+N$ $\stackrel{{\scriptstyle<}}{{\longrightarrow}}$ $B+A\;$ $\triangleright\ \text{initialisation}$ $(2)\;$ $A+B$ $\stackrel{{\scriptstyle<}}{{\longrightarrow}}$ $B+A\;$ $\triangleright\ \text{fix order}$ $(3)\;$ $A+N$ $\stackrel{{\scriptstyle<}}{{\longrightarrow}}$ $N+A\;$ $\triangleright\ \text{fix order}$ $(4)\;$ $N+B$ $\stackrel{{\scriptstyle<}}{{\longrightarrow}}$ $B+N\;$ $\triangleright\ \text{fix order}$
@@ 3.2 Fast Median Computation
2:Select randomly leader (as pivot) $p\in S$ $\triangleright$ leader election
@@
9:  $p\leftarrow$ randomly chosen (by current $p$) agent in $C$ $\triangleright$ leader hand over
@@
11:return $p$ $\triangleright$ result announcement
@@ 3.2.1 Partitioning via Coloring
If, after drawing i1i-1 balls, no more than (i1)p(i-1)p of them are black, we have $q_{i}=\Pr(X_{i}=1|\text{current urn content})\geq p$.
@@
  • If, after drawing i1i-1 balls, more than (i1)p(i-1)p of them are black, we independently draw $Y_{i}=1$ with a probability of $p$, and $Y_{i}=0$ otherwise.
  • If, after drawing i1i-1 balls, at most (i1)p(i-1)p of them are black and $X_{i}=0$, then $Y_{i}=0$.
  • If, after drawing i1i-1 balls, at most (i1)p(i-1)p of them are black and $X_{i}=1$, then $Y_{i}=1$ with a probability of $p/q_{i}$, and $Y_{i}=0$ otherwise.
Variables $Y_{1},Y_{2},\ldots,Y_{\kappa}$ are independent, because no $Y_{i}$ depends on any combination of $Y_{j}:j<i$, and $\Pr(Y_{i})=p$. If $X<(1-\delta)p{\kappa}$, then let $j\in[0,{\kappa})$ be the last draw before which not less than (j1)p(j-1)p balls drawn are black. This implies $Y_{i}\leq X_{i}$ for $i>j$ and $X_{1}+\cdots+X_{j}\geq p(j-1)$. Thus when $X_{1}+\cdots+X_{\kappa}<(1-\delta)p{\kappa}-p$, we get
@@
We divide the sequence of interactions into interaction periods $t_{0},t_{1}),[t_{1},t_{2}),[t_{2},t_{3}),\dots$ such that the first interaction period has length $O(n\log^{2}n),$ and each subsequent period is of length $O(n\log n)$. We show that in interaction $t_{\log n}$ all $m$ agents are in the group $\mathcal{G}$ whp. Each colored agent can utilise at most 21 tickets to move other agents to $\mathcal{G}$. This implies that in $t_{\log n}$ at least $m/22$ of these agents have color. Since each interaction between groups removes an agent from ${\mathcal{G}}_{in}$ irreversibly, and there are $O(n\log^{2}n)$ interactions altogether whp, by Lemma 2 the fragmented time of one phase is $O(\log^{2}n)$. The thesis of the lemma is a direct consequence of Lemma 9, which we formulate below. And indeed, since this implies $k(t_{\log n})\geq\min\left\{20\cdot 2^{\log n-1}\log^{2}n,m\right\},$ we get $k(t_{\log n})=m$. ∎
@@
We first deal with the base case $i=1$. Let $t=t_{0}+cn\log^{2}n$, where a constant $c>0$ is to be specified later on, and $t_{1}=t+500n$. Note that the pivot always colors an uncolored agent during an interaction. During interaction period $t_{0},t)$ the pivot is chosen as the initiator at least $\log^{2}n$ times whp. Indeed, by Chernoff bound the number of times $X,$ for which the pivot is chosen to be an initiator, satisfies for some $c>0$
@@
If in $t_{0},t)$ group ${\mathcal{G}}_{in}$ runs out of agents, then $k(t)=m$. Otherwise in time $t$ at least $\log^{2}n$ agents are colored and by lemma 7 we have $k(t_{1})=k(t+500n)\geq\min\left\{20\log^{2}n,m\right\}$ whp.
On the other hand, consider an uncolored agent from the first bucket. No matter whether it belongs to ${\mathcal{G}}_{in}$ or $\mathcal{G}$}, it has probability at least $\frac{1}{nm}$ of interaction with the pivot. Due to the way the buckets are defined, see the proof of Lemma 8, each of them has at least $\frac{m}{4\log n}$ agents. If in any interaction of $t_{0},t)$ more than $\frac{m}{10\log n}$ agents are colored, then $Inf_{i}(t_{1})>\min\left\{\log n,\frac{m}{10\log n}\right\}$. Otherwise, the probability of coloring a new agent from the first bucket in each interaction of $t_{0},t)$ is at least $\frac{3}{20n\log n}$. By Chernoff bound, for some $c>0$, the number $X$ of colored agents in the first bucket in $t$ satisfies
@@
then at least $y$ agents from the $i$th bucket are colored in $t_{i-1},t)$ or in some $\tau$ the number of uncolored agents in $\tau$ is smaller than $\frac{0.9k(\tau)}{8n\log n}$. Both cases imply